OpenStudy (anonymous):

Which shows ^6 sqr rt (24)^2 in simplified radical form with the smallest possible index?

OpenStudy (tanner23456):

Is this what you mean? \[\sqrt[6]{24^2}\] If that's the case put it in exponential form first. 24^(2/6) reduce the fraction in the exponent to simplest form and put it back in radical form

OpenStudy (anonymous):

yesh!! thanks the multiple choices r.. \[2 ^3\sqrt{3}\] \[3 \sqrt[3]{4}\] \[4 \sqrt[3]{3}\] \[6 \sqrt[3]{4}\]

OpenStudy (tanner23456):

ok so when you reduce the fraction you get \[24^{1/3} or \sqrt[3]{24}\] now factor out the 24 to 2*2*2*3 notice their are 3 2's so we can pull out a 2 so you're left with \[2\sqrt[3]{3}\]

OpenStudy (anonymous):

thank yoooooouuu :D *bug hug*

OpenStudy (anonymous):

the fuq? who r u?

OpenStudy (tanner23456):

Don't worry about it. @tywan has been doing this to several people.

OpenStudy (anonymous):

okay i was like gasp!! ive never done tht stuff